OpenStudy (anonymous):

Rationalize the denominator: 3 over √2.

OpenStudy (anonymous):

multiply sqrt2 by denominator and numerator. You get.....

OpenStudy (anonymous):

3sqrt2/2

OpenStudy (unklerhaukus):

\[\frac 3{\sqrt 2}=\frac 3{\sqrt 2}\times1=\frac 3{\sqrt 2}\times\frac {\sqrt 2}{\sqrt 2}=\]

OpenStudy (anonymous):